OpenStudy (anonymous):

Having some trouble factoring. I think I'm using the wrong factoring method, or that it can't be factored. The function I need to factor is f(t) = -16t^2 - 32t + 384.

OpenStudy (anonymous):

f(t) = -16t^2 - 32t + 384 = -16(t^2 +2t -24) can you now

OpenStudy (anonymous):

Would this be the correct first step to that? -16(t^2 + 2t - 24).

OpenStudy (anonymous):

Oh. So that is the correct first step?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

= -16(t^2 +2t -24) = = -16(t^2 +6t -4t -24)

OpenStudy (anonymous):

It's supposed to be like that, next? It IS asking me to "Factor the function f(t) and use the factors to interpret the meaning of the x-intercept of the function," but is that really how it is then? If so, what does it mean, exactly?

OpenStudy (anonymous):

@matricked Where in the world did you get that -4t?

OpenStudy (anonymous):

-24 = 6 x-4

OpenStudy (anonymous):

@matricked, so where is the t?

OpenStudy (anonymous):

If you're trying to get negative twenty-four, why still leave it in the expression?

OpenStudy (anonymous):

= -16(t^2 +2t -24) =-16 (t^2 +(6-4)t -24)

OpenStudy (anonymous):

Oh, I see. Your parens are messed up. You mean \[-16(t-4)(t+6)\] For the factored product.

OpenStudy (anonymous):

...That would make a lot more sense.

OpenStudy (anonymous):

@FrostFelon Yeah, I've never seen anyone factor that way. I makes sense if you were used to it. He is basically solving the problem, but instead of factoring, inserts the two terms where they are multiplied.

OpenStudy (anonymous):

Then factors simplifies I assume.

OpenStudy (anonymous):

=-16 (t^2 +(6-4)t -24) = -16 ( t^2 +6t -4t -24) =-16( t(t+6) -4(t+6)) =-16(t+6)(t-4)

OpenStudy (anonymous):

That actually takes longer than the leap of logic I usually use.

OpenStudy (anonymous):

That is super messed up. If it works for you, fine I guess, but you have not written what you think you have. According to your last step you have \[f(x)=-16(t^{2}+6t)+(-4t-24)\]

OpenStudy (anonymous):

Last step before you give -16(t+6)(t-4).

OpenStudy (anonymous):

Opps sorry; \[f(x)=-16((t^{2}+6t)+(-4t-24)\] So, \[f(x)=-16t^{2}+-96t+48t+384\] \[f(x)=-16t^{2}-48t+384\]

OpenStudy (anonymous):

There's another part of the problem that I'm not entirely sure about.

OpenStudy (anonymous):

Complete the square of the expression for f(x) to determine the vertex of the graph of f(x). Would this be a maximum or minimum on the graph? I can easily tell if it's a maximum or a minimum, but I don't know how to complete the square.

OpenStudy (anonymous):

The vertex is (-1, 400), right?

OpenStudy (anonymous):

@FrostFelon, Little late, but I believe complete the square means find t and solve \[16t^{2}\]

OpenStudy (anonymous):

Oh.